In a flag ceremony, the number of students in each row is 6 fewer than the number of rows. If there are 891 students in the flag ceremony, how many are in each row?
Answer by ankor@dixie-net.com(22740)

In a flag ceremony, the number of students in each row is 6 fewer than the number of rows.
If there are 891 students in the flag ceremony, how many are in each row?
:
Let x = no. of students
then
(x+6) = no. of rows
:
x(x+6) = 891
a quadratic equation
x^2 + 6x - 891 = 0
you can use the quadratic formula, but this will factor to:
(x+33) (x-27) = 0
the positive solution
x = 27 students in each row
:
:
Check this out (33 rows): 27 * 33 = 891